### 12 Feb 2024 v3

# Universality of almost periodicity in bounded discrete time series

Abstract.
We consider bounded discrete time series. From its statistical feature, without any use of the Fourier transform, we find a suitable almost periodic function which approximates the corresponding time series in a local time interval.
Key words and phrases:
almost periodic function, chaos in discrete time, recurrent neural network
2020 Mathematics Subject Classification:
Primary 42A20; Secondary 11B50
1. Introduction
In this paper we consider universality of almost periodicity. More precisely we show that, without any use of the Fourier transform, we find a suitable almost periodic function which approximates the corresponding bounded discrete time series in a local time interval.
Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}$ such that (for a technical reason, we employ transcendental numbers)
  • $\{a_{1}^{K},a_{2}^{K},\cdots,a^{K}_{K-1},a_{K}^{K}\}\subset\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\subset\mathbb{R}\setminus\mathbb{Q}$,
  • $-1=a_{0}^{K}<a_{1}^{K}<a_{2}^{K}<\cdots<a_{K+1}^{K}=1$,
  • $\sup_{1\leq k\leq K+1}|a_{k-1}^{K}-a_{k}^{K}|\leq^{\exists}\!\!C/K$.
We discretize $y$ as follows:
$\bar{y}^{K}(t):=\operatorname*{arg\,min}_{a\in\{a^{K}_{k}\}_{k=1}^{K}}|y(t)-(a-0)|,$
where $a-0:=a-\varepsilon$ for any sufficiently small $\varepsilon>0$. In what follows we sometimes omit the index $K$ unless confusion occurs. To define a suitable statistical feature, we need to classify patterns in $\bar{y}$. Let $\sigma_{n}$ ($n=1,2,\cdots,N$) be a permutation operator, namely,
$\sigma_{n}:\{1,2,\cdots,L\}\to\{a_{1},a_{2},\cdots,a_{K-1},a_{K}\}$
($\ell\mapsto\sigma_{n}(\ell)$) with $\sigma_{n}\not=\sigma_{n^{\prime}}$ ($n\not=n^{\prime}$), and we impose the following two conditions for determining $N$:
(1) $\begin{cases}\text{For any $t\in\mathbb{Z}$, there is $n\in\{1,\cdots,N\}$ such that}\\ \qquad\text{$\sigma_{n}(\ell)=\bar{y}(t-\ell)$ for $\ell=1,2,\cdots,L$},\\ \text{For any $n\in\{1,\cdots,N\}$ there is $t\in\mathbb{Z}$ such that}\\ \qquad\text{$\sigma_{n}(\ell)=\bar{y}(t-\ell)$ for $\ell=1,2,\cdots,L$.}\end{cases}$
Note that $N\leq K^{L}$ due to the sequence with repetition. The crucial statistical feature is the following: for any $y:\mathbb{Z}\to[-1,1]$, we have the following conditional probability: there exists $\delta^{K}\in[0,1)$ (depending on $K$) such that for any $n\in\{1,2,\cdots,N\}$,
(2) $P\bigg{(}a_{k(n)}=\bar{y}(\cdot)\bigg{|}\sigma_{n}(\ell)=y(\cdot-\ell)\quad\text{for}\quad\ell=1,2,\cdots,L\bigg{)}\geq 1-\delta^{K}$
with appropriately chosen $a_{k(n)}$ satisfying the following property:
  • Magnitude relation: For any $K^{\prime}>0$ and $t\in\mathbb{Z}$,
    (3) $\left|a^{K}_{k(n_{t}^{K})}-a^{K+K^{\prime}}_{k(n_{t}^{K+K^{\prime}})}\right|<\frac{C}{K},$
    where $n_{t}^{K}$ is determined as
    (4) $\bar{y}^{K}(t-\ell)=\sigma_{n_{t}^{K}}(\ell)\quad\text{for}\quad\ell=1,2,\cdots,L.$
Then we can state the main theorem as follows:
Theorem 1.
Assume $(\sigma_{n},a_{k(n)})$ satisfies the following:
  • (Uniqueness in backward direction.) For any $n^{\prime}\in\{1,2,\cdots,N\}$, there exists at most one $n\in\{1,2,\cdots,N\}$ such that
    (5) $(a_{k(n)},\sigma_{n}(1),\cdots,\sigma_{n}(L-1))=(\sigma_{n^{\prime}}(1),\sigma_{n^{\prime}}(2),\cdots,\sigma_{n^{\prime}}(L)).$
Then, for any $t_{0}\in\mathbb{Z}$, there exists an almost periodic function $u$ such that
(6) $\prod_{t^{\prime}=1}^{t}P\bigg{(}|u(t_{0}+t^{\prime})-y(t_{0}+t^{\prime})|<\frac{C}{K}\bigg{)}\geq\left(1-\delta^{K}\right)^{t}$
for any $t\in\mathbb{Z}_{\geq 1}$.
Remark 1.
Deriving the inequality (6) itself is rather obvious and quite natural due to (2). The nontrivial discovery is that $u$ is a discretized almost periodic function provided by (5), without any use of Fourier transform.
2. Proof of main theorem
Let $\sigma^{*}_{n}$ ($n=1,2,\cdots,N$) be an adjoint type of permutation operator, more precisely, let
(7) $\sigma^{*}_{n}(\ell-1):=\frac{1}{\sigma_{n}(\ell)}$
for $\ell\in\{1,\cdots,L\}$ and $n\in\{1,\cdots N\}$, where $\sigma^{*}_{i}(0):=\sigma_{i}^{*}(L)$. The definition of the following $N\times N$ matrix $X$ is the key to prove the main theorem:
$X:=\left[h\left(\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\right)\right]_{i,j}$
for some $h:\mathbb{R}\to[-1,1]$,
Lemma 2.
$X$ is a regular matrix provided by appropriately chosen $h:\mathbb{R}\to[-1,1]$.
Proof.
First, by the Lindemann–Weierstrass theorem, we see that
(8) $\frac{a_{k^{\prime}}}{a_{k}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-1\}\quad(k\not=k^{\prime}),$
(9) $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}}{a_{k_{\ell}}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-L,-L+1,\cdots,-1,0,1,\cdots,L-1\}$
except for the $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$ case. Also
$\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}=L$
if and only if $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$. By (7), (8), (9) and Lindemann–Weierstrass theorem, we see
$\sum_{\ell=1}^{L}\sigma^{*}_{j}(\ell-1)\sigma_{j}(\ell)=L\quad\text{and}\quad\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\not=L\quad(i\not=j).$
In order to construct an appropriate $h$, we use a finite set $G$ as follows:
$G:=\left\{\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell):i,j\in\{1,2,\cdots,N\}\right\}\subset\mathbb{R}.$
Note that $L\in G$. Now we take a smooth function $h:\mathbb{R}\to[-1,1]$ satisfying the following:
$\begin{split}&h(L)\in\mathbb{Q}\setminus\{0\},\\ &h(\gamma)\in\left(\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\cup\{0\}\right)\subset(\mathbb{R}\setminus\mathbb{Q})\cup\{0\},\quad\text{for}\quad\gamma\subset G\setminus\{1\}.\end{split}$
Then we can easily check that (applying the Lindemann–Weierstrass theorem)
  • $h(\gamma_{1})h(\gamma_{2})\cdots h(\gamma_{N})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}$  for  $\{\gamma_{n}\}_{n=1}^{N}\in G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
  • for any $\{\tau_{n^{\prime}}\}_{n^{\prime}=1}^{N!-1}\in\{-1,1\}^{N!-1}$ and $\{\{\gamma_{n,n^{\prime}}\}_{n=1}^{N}\}_{n^{\prime}=1}^{N!-1}\subset G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
    $\sum_{n^{\prime}=1}^{N!-1}\tau_{n^{\prime}}h(\gamma_{1,n^{\prime}})h(\gamma_{2,n^{\prime}})\cdots h(\gamma_{N,n^{\prime}})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
By applying the above properties, then we see that the determinant of the matrix $X$ is nonzero, since it is expressed as
$|X|=\eta_{1}+\eta_{2},\quad\eta_{1}\in\mathbb{Q}\setminus\{0\},\quad\eta_{2}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
Now we resume the proof of the main theorem. Let us define $N\times N$ matrices: $Y$ and $W$ as follows:
(10) $Y:=\left[\sum_{\ell=1}^{L}\sigma_{i}^{*}(\ell)\sigma_{j}(\ell)\right]_{i,j}\quad\text{and}\quad W:=YX^{-1}.$
By (1), for each $t\in\mathbb{Z}$, there exists a unique column vector
$e_{n_{t}}:=(\underbrace{\overbrace{0,0,\cdots,0,1}^{n_{t}},0\cdots,0}_{N})^{T}$
such that
(11) $\bar{y}(t-\ell)=\sigma(\ell)e_{n_{t}}\quad\text{for}\quad\ell=1,2,\cdots,L,$
where
$\sigma(\ell):=(\sigma_{1}(\ell),\sigma_{2}(\ell),\cdots,\sigma_{N}(\ell)).$
Thus we have
$W\left(h\left(\sum_{\ell=1}^{L}\sigma^{*}(\ell-1)\bar{y}(t-\ell)\right)\right)=WXe_{n_{t}}=Ye_{n_{t}}=\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t-\ell),$
where $\sigma^{*}(\ell):=(\sigma^{*}_{1}(\ell),\sigma^{*}_{2}(\ell),\cdots,\sigma_{N}^{*}(\ell))^{T}$. To find an appropriate almost periodic function $u$, we create the explicit formulas $\{\bar{u}(t)\}_{t\geq t_{0}}$ inductively from initial data $\{\bar{y}(t_{0}-\ell)\}_{\ell=1}^{L}$:
(12) $\begin{split}&\begin{cases}\bar{r}(t_{0})&=h(\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t_{0}-\ell)),\\ \bar{u}(t_{0})&=W^{out}\bar{r}(t_{0}),\end{cases}\\ &\begin{cases}\bar{r}(t)&=h(W\bar{r}(t-1)+\sigma^{*}(L)\bar{u}(t-1)-\sigma^{*}(L)\bar{u}(t-L-1)),\\ \bar{u}(t)&=W^{out}\bar{r}(t),\end{cases}\end{split}$
for $t=t_{0}+1,t_{0}+2,\cdots$, where $W^{out}:=AX^{-1}$ and $A:=(a_{k(1)},a_{k(2)},\cdots,a_{k(N)})$. Rigorously we regard $\bar{u}(t-L-1)$ as $\bar{y}(t-L-1)$ for $t=t_{0}+1,t_{0}+2\cdots,t_{0}+L$. Now we show that $\bar{u}^{K}(t)$ is a discretized periodic function for any $K$. By (2), we have a chain of the permutation operator $(\sigma^{K}_{j},a^{K}_{k(j)})$ as follows:
$(\sigma_{j^{K}_{1}},a^{K}_{k(j_{1})})\rightarrow(\sigma_{j^{K}_{2}},a^{K}_{k(j_{2})})\rightarrow(\sigma_{j^{K}_{3}},a^{K}_{k(j_{3})})\rightarrow\cdots.$
Note that the sequence $j_{1},j_{2},j_{3}\cdots$ is different from $n_{1},n_{2},n_{3}\cdots$ in (3). Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite and by backward uniqueness (5), then there is $m\in\mathbb{Z}_{\geq 2}$ such that $(\sigma_{j^{K}_{1}}a_{k(j_{1})}^{K})=(\sigma_{j^{K}_{m}},a_{k(j_{m})}^{K})$. This means that we have a periodic chain for any $K$. Moreover $\lim_{K\to\infty}\bar{u}^{K}=:u$ is uniform convergence due to (3). Thus $u$ is almost periodic (see Corduneanu [1, (1.2)]).
Remark 2.
In some situations, this periodic function $\bar{u}^{K}$ may be different for different choice of $t_{0}$, but we do not go into the detail here.
Acknowledgments. The author is grateful to Professors Chikara Nakayama and Yoshitaka Saiki for valuable comments. Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 20H01819.

### 01 Mar 2024 v4

@@
 Abstract.
-We consider bounded discrete time series.
-From its statistical feature, without any use of the Fourier transform, we find a suitable almost periodic function which approximates the corresponding time series in a local time interval.
+We consider arbitrary bounded discrete time series.
+From its statistical feature, without any use of the Fourier transform, we find an almost periodic function which suitably characterizes the corresponding time series.
 Key words and phrases:
@@
 In this paper we consider universality of almost periodicity.
-More precisely we show that, without any use of the Fourier transform, we find a suitable almost periodic function which approximates the corresponding bounded discrete time series in a local time interval.
-Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}$ such that (for a technical reason, we employ transcendental numbers)
+More precisely, without any use of the Fourier transform, we construct a suitable recursive formula which approximates the corresponding bounded discrete time series in a local time interval, and the corresponding explicit formula tends to an almost periodic function as time tends to infinity.
+Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series arbitrarily chosen, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}$ such that (for a technical reason, we employ transcendental numbers)
   • $\{a_{1}^{K},a_{2}^{K},\cdots,a^{K}_{K-1},a_{K}^{K}\}\subset\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\subset\mathbb{R}\setminus\mathbb{Q}$,
@@
 We discretize $y$ as follows:
-$\bar{y}^{K}(t):=\operatorname*{arg\,min}_{a\in\{a^{K}_{k}\}_{k=1}^{K}}|y(t)-(a-0)|,$
+(1) $\bar{y}_{K}(t):=\operatorname*{arg\,min}_{a\in\{a^{K}_{k}\}_{k=1}^{K}}|y(t)-(a-0)|,$
 where $a-0:=a-\varepsilon$ for any sufficiently small $\varepsilon>0$.
@@
 Note that $N\leq K^{L}$ due to the sequence with repetition.
-The crucial statistical feature is the following: for any $y:\mathbb{Z}\to[-1,1]$, we have the following conditional probability: there exists $\delta^{K}\in[0,1)$ (depending on $K$) such that for any $n\in\{1,2,\cdots,N\}$,
-(2) $P\bigg{(}a_{k(n)}=\bar{y}(\cdot)\bigg{|}\sigma_{n}(\ell)=y(\cdot-\ell)\quad\text{for}\quad\ell=1,2,\cdots,L\bigg{)}\geq 1-\delta^{K}$
+The crucial statistical feature is the following: for any $y:\mathbb{Z}\to[-1,1]$, we have the following conditional probability: there exists a $\delta_{K}\in[0,1)$ (of course depending on $K$) such that for any $n\in\{1,2,\cdots,N\}$,
+(3) $P\bigg{(}a_{k(n)}=\bar{y}(\cdot)\bigg{|}\sigma_{n}(\ell)=\bar{y}(\cdot-\ell)\quad\text{for}\quad\ell=1,2,\cdots,L\bigg{)}\geq 1-\delta_{K}$
 with appropriately chosen $a_{k(n)}$ satisfying the following property:
@@
     (4) $\left|a^{K}_{k(n_{t}^{K})}-a^{K+K^{\prime}}_{k(n_{t}^{K+K^{\prime}})}\right|<\frac{C}{K},$
-    where $n_{t}^{K}$ is determined as
-    (4) $\bar{y}^{K}(t-\ell)=\sigma_{n_{t}^{K}}(\ell)\quad\text{for}\quad\ell=1,2,\cdots,L.$
+    where $n_{t}^{K}\in\{1,2,\cdots,N\}$ is determined as
+    (5) $\bar{y}_{K}(t-\ell)=\sigma_{n_{t}^{K}}(\ell)\quad\text{for}\quad\ell=1,2,\cdots,L.$
 Then we can state the main theorem as follows:
 Theorem 1.
-Assume $(\sigma_{n},a_{k(n)})$ satisfies the following:
-  • (Uniqueness in backward direction.) For any $n^{\prime}\in\{1,2,\cdots,N\}$, there exists at most one $n\in\{1,2,\cdots,N\}$ such that
-    (5) $(a_{k(n)},\sigma_{n}(1),\cdots,\sigma_{n}(L-1))=(\sigma_{n^{\prime}}(1),\sigma_{n^{\prime}}(2),\cdots,\sigma_{n^{\prime}}(L)).$
-Then, for any $t_{0}\in\mathbb{Z}$, there exists an almost periodic function $u$ such that
-(6) $\prod_{t^{\prime}=1}^{t}P\bigg{(}|u(t_{0}+t^{\prime})-y(t_{0}+t^{\prime})|<\frac{C}{K}\bigg{)}\geq\left(1-\delta^{K}\right)^{t}$
-for any $t\in\mathbb{Z}_{\geq 1}$.
+For any $t_{0}\in\mathbb{Z}$ there exists $u:\mathbb{Z}\to[-1,1]$ such that the following inequality holds:
+(6) $\prod_{t^{\prime}=0}^{t}P\bigg{(}|u(t_{0}+t^{\prime})-y(t_{0}+t^{\prime})|<\frac{C}{K}\bigg{)}\geq\left(1-\delta_{K}\right)^{t}$
+for any $K\in\mathbb{Z}_{\geq 1}$ and $t\in\mathbb{Z}_{\geq 0}$. Moreover there exists a discretized almost periodic function $ap:\mathbb{Z}\to[-1,1]$ such that
+[…]
 Remark 1.
 Deriving the inequality (6) itself is rather obvious and quite natural due to (3).
-The nontrivial discovery is that $u$ is a discretized almost periodic function provided by (5), without any use of Fourier transform.
+The nontrivial discovery is explicit construction of $u$ (see (12)), and it tends to the discretized almost periodic function, without any use of Fourier transform.
+Corollary 2.
+Let us mention the dynamical system case (c.f. [1, 2]). With the aid of Takens’ embedding theorem [6] (see also [5]), it suffices to consider in a delay coordinate. Let
+$\begin{split}Y_{t}&:=(y(t),y(t-1),\cdots,y(t-L+1))^{T},\\ \overline{Y}_{t}&:=(\bar{y}(t),\bar{y}(t-1),\cdots,\bar{y}(t-L+1))^{T},\\ \overline{U}_{t}&:=(\bar{u}(t),\bar{u}(t-1),\cdots,\bar{u}(t-L+1))^{T},\end{split}$
+where $\bar{u}$ is generated by (12), and let $\Phi:[-1,1]^{L}\to[-1,1]^{L}$ be such that $Y_{t-1}\mapsto Y_{t}$ (in Appendix 3, we give the simplest example of $\Phi$).
+We see that, for any $t\in\mathbb{Z}_{\geq 0}$, there exists $t^{*}\in\mathbb{Z}_{\geq 0}$ such that
+$\begin{cases}\displaystyle\overline{Y}_{t_{0}+t^{*}}=\overline{U}_{t_{0}+t},\quad|Y_{t_{0}+t^{*}}-\overline{U}_{t_{0}+t}|\leq\frac{\sqrt{L}C}{K},\\ Y_{t_{0}+t^{*}}=\Phi(Y_{t_{0}+t^{*}-1}),\\ \displaystyle\overline{Y}_{t_{0}+t^{*}-1}=\overline{U}_{t_{0}+t-1},\quad|Y_{t_{0}+t^{*}-1}-\overline{U}_{t_{0}+t-1}|\leq\frac{\sqrt{L}C}{K}.\end{cases}$
+Let
+$\beta:=\sup_{t\not=t^{\prime}}|\frac{|\Phi(Y(t))-\Phi(Y(t^{\prime}))|}{|Y(t)-Y(t^{\prime})|}$
+and assume $\beta\in\mathbb{R}_{\geq 1}$. By applying the above inequalities inductively, then we have
+$\begin{split}&|\bar{u}_{K}(t_{0}+t)-y(t_{0}+t)|\leq|\overline{U}(t_{0}+t)-Y(t_{0}+t)|\\ \leq&|Y(t_{0}+t^{*})-Y(t_{0}+t)|+\frac{\sqrt{L}C}{K}\\ \leq&\beta|Y(t_{0}+t^{*}-1)-Y(t_{0}+t-1)|+\frac{\sqrt{L}C}{K}\\ \leq&\beta|\overline{U}(t_{0}+t-1)-Y(t_{0}+t-1)|+\beta\frac{\sqrt{L}C}{K}+\frac{\sqrt{L}C}{K}\leq\cdots\\ \lesssim&\beta^{t}\left|\overline{U}\left(t_{0}\right)-Y\left(t_{0}\right)\right|+\beta^{t}\frac{\sqrt{L}C}{K}\\ \lesssim&\frac{\sqrt{L}C\beta^{t}}{K}.\end{split}$
+This means that
+$P\bigg{(}|\bar{u}_{K}(t_{0}+t)-y(t_{0}+t)|\lesssim\frac{\sqrt{L}C\beta^{t}}{K}\bigg{)}=1,$
+that is,
+[…]
+Note that the case $\beta<1$ is trivial, since $\Phi$ becomes a contraction mapping. Again, there exists a discretized almost periodic function $ap:\mathbb{Z}\to[-1,1]$ such that
+$|u(t)-ap(t)|\to 0\quad(t\to\infty).$
 2. Proof of main theorem
@@
 where $\sigma^{*}(\ell):=(\sigma^{*}_{1}(\ell),\sigma^{*}_{2}(\ell),\cdots,\sigma_{N}^{*}(\ell))^{T}$.
-To find an appropriate almost periodic function $u$, we create the explicit formulas $\{\bar{u}(t)\}_{t\geq t_{0}}$ inductively from initial data $\{\bar{y}(t_{0}-\ell)\}_{\ell=1}^{L}$:
+To find an appropriate function $u$, we create the explicit formulas $\{\bar{u}(t)\}_{t\geq t_{0}}$ inductively from initial data $\{\bar{y}(t_{0}-\ell)\}_{\ell=1}^{L}$:
 (12) $\begin{split}&\begin{cases}\bar{r}(t_{0})&=h(\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t_{0}-\ell)),\\ \bar{u}(t_{0})&=W^{out}\bar{r}(t_{0}),\end{cases}\\ &\begin{cases}\bar{r}(t)&=h(W\bar{r}(t-1)+\sigma^{*}(L)\bar{u}(t-1)-\sigma^{*}(L)\bar{u}(t-L-1)),\\ \bar{u}(t)&=W^{out}\bar{r}(t),\end{cases}\end{split}$
@@
 Rigorously we regard $\bar{u}(t-L-1)$ as $\bar{y}(t-L-1)$ for $t=t_{0}+1,t_{0}+2\cdots,t_{0}+L$.
-Now we show that $\bar{u}^{K}(t)$ is a discretized periodic function for any $K$.
-By (2), we have a chain of the permutation operator $(\sigma^{K}_{j},a^{K}_{k(j)})$ as follows:
-$(\sigma_{j^{K}_{1}},a^{K}_{k(j_{1})})\rightarrow(\sigma_{j^{K}_{2}},a^{K}_{k(j_{2})})\rightarrow(\sigma_{j^{K}_{3}},a^{K}_{k(j_{3})})\rightarrow\cdots.$
+Now we show that $\bar{u}_{K}(t)$ tends to a discretized periodic function for any $K$.
+By (3), for each $K$, we have a chain of the permutation operator $(\sigma_{j},a_{k(j)})$ as follows:
+$(\sigma_{j_{1}},a_{k(j_{1})})\rightarrow(\sigma_{j_{2}},a_{k(j_{2})})\rightarrow(\sigma_{j_{3}},a_{k(j_{3})})\rightarrow\cdots.$
 Note that the sequence $j_{1},j_{2},j_{3}\cdots$ is different from $n_{1},n_{2},n_{3}\cdots$ in (4).
-Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite and by backward uniqueness (5), then there is $m\in\mathbb{Z}_{\geq 2}$ such that $(\sigma_{j^{K}_{1}}a_{k(j_{1})}^{K})=(\sigma_{j^{K}_{m}},a_{k(j_{m})}^{K})$.
-This means that we have a periodic chain for any $K$.
-Moreover $\lim_{K\to\infty}\bar{u}^{K}=:u$ is uniform convergence due to (3).
-Thus $u$ is almost periodic (see Corduneanu [1, (1.2)]).
+Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite, then there exist $m,M\in\mathbb{Z}_{\geq 2}$ ($m<M$) such that $(\sigma_{j_{m}}a_{k(j_{m})})=(\sigma_{j_{M}},a_{k(j_{M})})$.
+This means that there exists $T_{K}$ such that $\bar{u}_{K}(t)$ is a periodic chain for $t>T_{K}$.
+Moreover $\lim_{K\to\infty}\bar{u}_{K}=:u$ is uniform convergence due to (4).
+Thus $u(t)$ tends to an almost periodic function for $t\to\infty$ (see Corduneanu [3, (1.2)]).
 Remark 2.
-In some situations, this periodic function $\bar{u}^{K}$ may be different for different choice of $t_{0}$, but we do not go into the detail here.
+In some situations, this periodic chain may be different for different choice of $t_{0}$, but we do not go into the detail here.
+3. Appendix
+In this appendix, we give the simplest example of the dynamical system, that is, a linear regression relation.
+Let $L\in 2\mathbb{Z}_{\geq 1}$.
+For $\{p_{\ell}\}_{\ell=1}^{L}\subset\mathbb{R}$, we consider the following:
+(13) $y(t)=p_{1}y(t-1)+p_{2}y(t-2)+\cdots+p_{L-1}y(t-L+1)+p_{L}y(t-L)\quad(t\geq 0)$
+with initial data $\{y(-\ell)\}_{\ell=1}^{L}$. First we figure out the corresponding characteristic equation, that is, we plug the following:
+$y(-\ell)=\mu^{L-\ell},\quad(\mu\in\mathbb{R},\ \ell=0,1,\cdots,L)$
+into (13). After factorization of the $L$-th degree polynomial, then, we see that $\{p_{\ell}\}_{\ell}$ satisfies the following equality (at least we can easily figure out that $p_{L}=-1$):
+$\prod_{j=1}^{L/2}(\mu-e^{i\lambda_{j}})(\mu-e^{-i\lambda_{j}})=\mu^{L}-\sum_{\ell=1}^{L}p_{\ell}\mu^{L-\ell}=0$
+if and only if $y$ is a quasi-periodic function:
+(14) $y(t)=\sum_{j=1}^{L/2}a_{j}\sin\left(\lambda_{j}(t-b_{j})\right),\quad t\in\mathbb{Z}.$
+for $\{a_{j}\}_{j=1}^{L/2}$, $\{b_{j}\}_{j=1}^{L/2}$, $\{\lambda_{j}\}_{j=1}^{L/2}\subset\mathbb{R}$. Otherwise, the time series is not bounded.
 Acknowledgments.
-The author is grateful to Professors Chikara Nakayama and Yoshitaka Saiki for valuable comments.
+The author is grateful to Professors Masanobu Inubushi, Chikara Nakayama and Yoshitaka Saiki for valuable comments.
 Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 20H01819.

### 25 Mar 2024 v5

@@
 Key words and phrases:
-almost periodic function, chaos in discrete time, recurrent neural network
+almost periodic function, chaos in discrete time, recurrent neural network, topological ring
 2020 Mathematics Subject Classification:
@@
 More precisely, without any use of the Fourier transform, we construct a suitable recursive formula which approximates the corresponding bounded discrete time series in a local time interval, and the corresponding explicit formula tends to an almost periodic function as time tends to infinity.
-Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series arbitrarily chosen, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}$ such that (for a technical reason, we employ transcendental numbers)
-  • $\{a_{1}^{K},a_{2}^{K},\cdots,a^{K}_{K-1},a_{K}^{K}\}\subset\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\subset\mathbb{R}\setminus\mathbb{Q}$,
+Let $y:\mathbb{Z}\to[-1,1]$ be a bounded discrete time series arbitrarily chosen, and first let us discretize the range $[-1,1]$ as follows: For $K\in\mathbb{Z}_{\geq 1}$, we choose $\{a_{k}^{K}\}_{k=0}^{K+1}\subset[-1,1]$ such that
   • $-1=a_{0}^{K}<a_{1}^{K}<a_{2}^{K}<\cdots<a_{K+1}^{K}=1$,
+  • $a_{i}^{K}\not=0$ for any $1\leq i\leq K$ and
+    $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}\not=L$ for each $(k_{1},\ldots,k_{L})\not=(k^{\prime}_{1},\ldots,k^{\prime}_{L})$,
   • $\sup_{1\leq k\leq K+1}|a_{k-1}^{K}-a_{k}^{K}|\leq^{\exists}\!\!C/K$.
-We discretize $y$ as follows:
+Note that the above second condition is almost always satisfied in the following sense.
+Proposition 1.
+There is a closed set $Z$ of $[-1,1]^{K}$ whose Lebesgue measure is zero such that the above second condition is satisfied as soon as $(a^{K}_{1},\ldots,a^{K}_{K})$ is not in $Z$.
+Proof.
+We drop the superscipt $K$ for simplicity.
+Regard $a_{i}$ ($1\leq i\leq K$) as indeterminates.
+Clearly the measure of the set where $a_{i}=0$ is zero.
+Under the assumption that any $a_{i}$ is not zero, the condition $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}}{a_{k_{\ell}}}\not=L$ is equivalent to $\sum_{\ell}a_{k_{1}}\cdots a_{k^{\prime}_{\ell}}\cdots a_{k_{L}}-L\prod_{\ell}a_{k_{\ell}}\not=0$, the left-hand-side of which is a nonzero polynomial whenever $k_{\ell}\not=k^{\prime}_{\ell}$ for some $\ell$ (then the term $\prod_{\ell}a_{k_{\ell}}$ does not vanish).
+Hence by the following simple fact, the inequality holds for almost all values of $a_{i}$:
+$\bullet$ Let $f$ be a nonzero polynomial in $M$ variables. Then the Lebesgue measure of the zero point set $\{x\in\mathbb{R}^{M}:f(x)=0\}$ is zero.
+We now discretize $y$ as follows:
 (1) $\bar{y}_{K}(t):=\operatorname*{arg\,min}_{a\in\{a^{K}_{k}\}_{k=1}^{K}}|y(t)-(a-0)|,$
 where $a-0:=a-\varepsilon$ for any sufficiently small $\varepsilon>0$.
-In what follows we sometimes omit the index $K$ unless confusion occurs.
+In what follows we sometimes omit the index $K$ unless confusion occurs (for example, $\bar{y}=\bar{y}_{K}$ and $\bar{u}=\bar{u}_{K}$).
 To define a suitable statistical feature, we need to classify patterns in $\bar{y}$.
@@
 $\begin{split}Y_{t}&:=(y(t),y(t-1),\cdots,y(t-L+1))^{T},\\ \overline{Y}_{t}&:=(\bar{y}(t),\bar{y}(t-1),\cdots,\bar{y}(t-L+1))^{T},\\ \overline{U}_{t}&:=(\bar{u}(t),\bar{u}(t-1),\cdots,\bar{u}(t-L+1))^{T},\end{split}$
-where $\bar{u}$ is generated by (12), and let $\Phi:[-1,1]^{L}\to[-1,1]^{L}$ be such that $Y_{t-1}\mapsto Y_{t}$ (in Appendix 3, we give the simplest example of $\Phi$).
-We see that, for any $t\in\mathbb{Z}_{\geq 0}$, there exists $t^{*}\in\mathbb{Z}_{\geq 0}$ such that
+where $\bar{u}$ is generated by (10), and assume there exists a map $\Phi:[-1,1]^{L}\to[-1,1]^{L}$ such that $Y_{t-1}\mapsto Y_{t}$ (in Appendix 4, we give the simplest example of $\Phi$).
+Then we see that, for any $t\in\mathbb{Z}_{\geq 0}$, there exists $t^{*}\in\mathbb{Z}_{\geq 0}$ such that
 $\begin{cases}\displaystyle\overline{Y}_{t_{0}+t^{*}}=\overline{U}_{t_{0}+t},\quad|Y_{t_{0}+t^{*}}-\overline{U}_{t_{0}+t}|\leq\frac{\sqrt{L}C}{K},\\ Y_{t_{0}+t^{*}}=\Phi(Y_{t_{0}+t^{*}-1}),\\ \displaystyle\overline{Y}_{t_{0}+t^{*}-1}=\overline{U}_{t_{0}+t-1},\quad|Y_{t_{0}+t^{*}-1}-\overline{U}_{t_{0}+t-1}|\leq\frac{\sqrt{L}C}{K}.\end{cases}$
@@
 $P\bigg{(}|\bar{u}_{K}(t_{0}+t)-y(t_{0}+t)|\lesssim\frac{\sqrt{L}C\beta^{t}}{K}\bigg{)}=1,$
-that is,
+in other words,
 […]
@@
 (7) $\sigma^{*}_{n}(\ell-1):=\frac{1}{\sigma_{n}(\ell)}$
-for $\ell\in\{1,\cdots,L\}$ and $n\in\{1,\cdots N\}$, where $\sigma^{*}_{i}(0):=\sigma_{i}^{*}(L)$. The definition of the following $N\times N$ matrix $X$ is the key to prove the main theorem:
+for $\ell\in\{1,\cdots,L\}$ and $n\in\{1,\cdots,N\}$, and $\sigma^{*}_{i}(L):=\sigma_{i}^{*}(0)$.
+The definition of the following $N\times N$ matrix $X$ and a finite set $G\subset\mathbb{R}$ are the keys to prove the main theorem:
 $X:=\left[h\left(\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\right)\right]_{i,j}$
 for some $h:\mathbb{R}\to[-1,1]$,
+$G:=\left\{\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell):i,j\in\{1,2,\cdots,N\}\right\}.$
+Note that $L\in G$.
 Lemma 3.
-$X$ is a regular matrix provided by appropriately chosen $h:\mathbb{R}\to[-1,1]$.
+$X$ is a regular matrix for almost all $h\colon\mathbb{R}\to[-1,1]$ in the following sense.
+Let $\gamma_{1},\ldots,\gamma_{M}$ be the all elements of $G$.
+Then there is a closed set $Z$ of $\mathbb{R}^{M}$ whose Lebesgue measure is zero such that $X$ is regular as soon as $(h(\gamma_{1}),\ldots,h(\gamma_{M}))$ is not in $Z$.
 Proof.
-First, by the Lindemann–Weierstrass theorem, we see that
-(8) $\frac{a_{k^{\prime}}}{a_{k}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-1\}\quad(k\not=k^{\prime}),$
-(9) $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}}{a_{k_{\ell}}}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{-L,-L+1,\cdots,-1,0,1,\cdots,L-1\}$
-except for the $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$ case. Also
-$\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}=L$
-if and only if $k_{1}=k^{\prime}_{1},\ k_{2}=k^{\prime}_{2},\cdots,k_{L}=k^{\prime}_{L}$. By (7), (8), (9) and Lindemann–Weierstrass theorem, we see
-$\sum_{\ell=1}^{L}\sigma^{*}_{j}(\ell-1)\sigma_{j}(\ell)=L\quad\text{and}\quad\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell)\not=L\quad(i\not=j).$
-In order to construct an appropriate $h$, we use a finite set $G$ as follows:
-$G:=\left\{\sum_{\ell=1}^{L}\sigma^{*}_{i}(\ell-1)\sigma_{j}(\ell):i,j\in\{1,2,\cdots,N\}\right\}\subset\mathbb{R}.$
-Note that $L\in G$. Now we take a smooth function $h:\mathbb{R}\to[-1,1]$ satisfying the following:
-$\begin{split}&h(L)\in\mathbb{Q}\setminus\{0\},\\ &h(\gamma)\in\left(\{\pm e^{-\frac{n}{m}}\ ;m,n\in\mathbb{Z}_{\geq 1}\}\cup\{0\}\right)\subset(\mathbb{R}\setminus\mathbb{Q})\cup\{0\},\quad\text{for}\quad\gamma\subset G\setminus\{1\}.\end{split}$
-Then we can easily check that (applying the Lindemann–Weierstrass theorem)
-  • $h(\gamma_{1})h(\gamma_{2})\cdots h(\gamma_{N})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}$  for  $\{\gamma_{n}\}_{n=1}^{N}\in G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
-  • for any $\{\tau_{n^{\prime}}\}_{n^{\prime}=1}^{N!-1}\in\{-1,1\}^{N!-1}$ and $\{\{\gamma_{n,n^{\prime}}\}_{n=1}^{N}\}_{n^{\prime}=1}^{N!-1}\subset G^{N}\setminus\{\underbrace{1,1,\cdots,1}_{N}\}$,
-    $\sum_{n^{\prime}=1}^{N!-1}\tau_{n^{\prime}}h(\gamma_{1,n^{\prime}})h(\gamma_{2,n^{\prime}})\cdots h(\gamma_{N,n^{\prime}})\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
-By applying the above properties, then we see that the determinant of the matrix $X$ is nonzero, since it is expressed as
-$|X|=\eta_{1}+\eta_{2},\quad\eta_{1}\in\mathbb{Q}\setminus\{0\},\quad\eta_{2}\in(\mathbb{R}\setminus\mathbb{Q})\cup\{0\}.$
-Now we resume the proof of the main theorem. Let us define $N\times N$ matrices: $Y$ and $W$ as follows:
-(10) $Y:=\left[\sum_{\ell=1}^{L}\sigma_{i}^{*}(\ell)\sigma_{j}(\ell)\right]_{i,j}\quad\text{and}\quad W:=YX^{-1}.$
+Take a bijection $x\colon\gamma_{m}\mapsto x_{m}$ from $G$ to the set of indeterminates $\{x_{1},\ldots,x_{M}\}$.
+Consider the determinant $D$ of the matrix $[x(\sum\sigma_{i}^{*}(\ell-1)\sigma_{j}(\ell))]_{i,j\in N}$ with polynomial coefficients.
+Then $D$ is a nonzero polynomial because it is a monic polynomial of degree $N$ with respect to the indeterminate $x(L)$ (the leading term $x(L)^{N}$ comes from the diagonal).
+Now we apply the simple fact in the proof of Proposition 1 to $D$.
+Let $Z$ be the zero point set of $D$.
+Assume that $(h(\gamma_{1}),\ldots,h(\gamma_{M}))$ is not in $Z$.
+Then $\det X$ defined by $h$ is not zero so that $X$ is regular.
+∎
+With the aid of the inverse of $X$, we can define a $N\times N$ matrix $W$ as follows:
+(8) $W:=YX^{-1}\quad\text{with}\quad Y:=\left[\sum_{\ell=1}^{L}\sigma_{i}^{*}(\ell)\sigma_{j}(\ell)\right]_{i,j}.$
+Remark 2.
+There is a smooth function $h\colon\mathbb{R}\to\mathbb{R}$ such that the absolute value of any eigenvalue of the $W$ and $W^{T}W$ are strictly less than $1$ (c.f.
+[3]).
+This can be shown as follows: Let $c$ be a real number which is strictly greater than the absolute value of any eignevalue of $Y$ and the square root of the absolute value of any eigenvalue of $Y^{T}Y$.
+Then we can take a smooth function $h$ whose support does not contain any $\sum_{\ell=1}^{L}\frac{a_{k^{\prime}_{\ell}}^{K}}{a_{k_{\ell}}^{K}}$ for $(k_{1},\ldots,k_{L})\not=(k^{\prime}_{1},\ldots,k^{\prime}_{L})$ such that $h(L)=c$.
+Then $X=cI$, where $I$ is the unit matrix, and $W=\frac{1}{c}Y$, $W^{T}W=c^{-2}Y^{T}Y$.
 By (2), for each $t\in\mathbb{Z}$, there exists a unique column vector
@@
 To find an appropriate function $u$, we create the explicit formulas $\{\bar{u}(t)\}_{t\geq t_{0}}$ inductively from initial data $\{\bar{y}(t_{0}-\ell)\}_{\ell=1}^{L}$:
-(12) $\begin{split}&\begin{cases}\bar{r}(t_{0})&=h(\sum_{\ell=1}^{L}\sigma^{*}(\ell)\bar{y}(t_{0}-\ell)),\\ \bar{u}(t_{0})&=W^{out}\bar{r}(t_{0}),\end{cases}\\ &\begin{cases}\bar{r}(t)&=h(W\bar{r}(t-1)+\sigma^{*}(L)\bar{u}(t-1)-\sigma^{*}(L)\bar{u}(t-L-1)),\\ \bar{u}(t)&=W^{out}\bar{r}(t),\end{cases}\end{split}$
+(10) $\begin{split}&\begin{cases}\bar{r}(t_{0})&:=h(\sum_{\ell=1}^{L}\sigma^{*}(\ell-1)\bar{y}(t_{0}-\ell)),\\ \bar{u}(t_{0})&:=W^{out}\bar{r}(t_{0}),\end{cases}\\ &\begin{cases}\bar{r}(t)&:=h(W\bar{r}(t-1)+\sigma^{*}(L)\bar{u}(t-1)-\sigma^{*}(L)\bar{u}(t-L-1)),\\ \bar{u}(t)&:=W^{out}\bar{r}(t),\end{cases}\end{split}$
 for $t=t_{0}+1,t_{0}+2,\cdots$, where $W^{out}:=AX^{-1}$ and $A:=(a_{k(1)},a_{k(2)},\cdots,a_{k(N)})$.
 Rigorously we regard $\bar{u}(t-L-1)$ as $\bar{y}(t-L-1)$ for $t=t_{0}+1,t_{0}+2\cdots,t_{0}+L$.
+Note that, since $\bar{r}(t)=Xe_{n_{t^{\prime}}}$ for some $t^{\prime}$ ($t^{\prime}$ is not necessarily equal to $t$), we see $\bar{u}(t)=W^{out}\bar{r}(t)=Ae_{n_{t^{\prime}}}$.
 Now we show that $\bar{u}_{K}(t)$ tends to a discretized periodic function for any $K$.
@@
 Note that the sequence $j_{1},j_{2},j_{3}\cdots$ is different from $n_{1},n_{2},n_{3}\cdots$ in (4).
-Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite, then there exist $m,M\in\mathbb{Z}_{\geq 2}$ ($m<M$) such that $(\sigma_{j_{m}}a_{k(j_{m})})=(\sigma_{j_{M}},a_{k(j_{M})})$.
+Since the number of permutation operators $\{\sigma_{n}\}_{n=1}^{N}$ is finite, then there exist $m,M\in\mathbb{Z}_{\geq 2}$ ($m<M$) such that $(\sigma_{j_{m}},a_{k(j_{m})})=(\sigma_{j_{M}},a_{k(j_{M})})$.
 This means that there exists $T_{K}$ such that $\bar{u}_{K}(t)$ is a periodic chain for $t>T_{K}$.
 Moreover $\lim_{K\to\infty}\bar{u}_{K}=:u$ is uniform convergence due to (4).
-Thus $u(t)$ tends to an almost periodic function for $t\to\infty$ (see Corduneanu [3, (1.2)]).
-Remark 2.
-In some situations, this periodic chain may be different for different choice of $t_{0}$, but we do not go into the detail here.
+Thus $u(t)$ tends to an almost periodic function for $t\to\infty$ (see Corduneanu [4, (1.2)]).
 4. Appendix
@@
 Acknowledgments.
-The author is grateful to Professors Masanobu Inubushi, Chikara Nakayama and Yoshitaka Saiki for valuable comments.
+The authors are grateful to Professors Masanobu Inubushi and Yoshitaka Saiki for valuable comments.
+Also the first author thanks J.
+C.
+for leading him to join this work.
 Research of TY was partly supported by the JSPS Grants-in-Aid for Scientific Research 20H01819.
+Research of CN was partly supported by the JSPS Grants-in-Aid for Scientific Research 21K03199.